# MDIT: A Model-free Data Interpolation Method for Diverse Instruction Tuning

### v1

@@ Abstract
Yangning Li22footnotemark: 2    Zihua Lan22footnotemark: 2    Lv Qingsong     Yinghui Li    Hai-Tao Zheng33footnotemark: 3 Tsinghua University
@@ 1 Introduction
Instruction tuning has enabled large language models (LLMs) to accurately follow human instructions and significantly enhance their performance Dong et al. (2023); Longpre et al. (2023); Zhang et al. (2023); Yi et al. (2024); Li et al. (2025a). The diversity of instruction datasets plays an essential role in improving LLM’s ability to handle various scenarios Li et al. (2022c, 2023c); Muscato et al. (2024); Fan et al. (2025). Therefore, recent research focuses on curating high-diversity and wide-ranging instruction datasets Mukherjee et al. (2023); Chung et al. (2024); Xu et al. (2025); Li et al. (2024b).
@@ 2.1.1 Instruction Data Selection
Data selection methods aim to filter out and remove low-diversity instruction data, including metric-based and model-based methods Li et al. (2022a, 2023a); Ye et al. (2023c); Li et al. (2023d); Chen et al. (2023b); Qiu et al. (2024). Metric-based selections Gonen et al. (2022); Zhou et al. (2023); Ye et al. (2023b); Li et al. (2023b); Zeng et al. (2025) use quantitative metrics to identify diverse instruction data. Instruction mining Cao et al. (2023) uses a linear equation to evaluate instruction quality, while InstructionGPT-4 Wei et al. (2023) further filters multimodal instruction data Li et al. (2024c); Kuang et al. (2024); Yu et al. (2024a); Chen et al. (2024a) according to CLIP scores Radford et al. (2021); Li et al. (2024f) and instruction length. Model-based selections Wu et al. (2023); Chen et al. (2023a); Yu et al. (2023); Ge et al. (2024) leverage LLMs as data selectors to identify more diverse instructions Li et al. (2024a); Liu et al. (2024). INSTAG Lu et al. (2023) utilizes ChatGPT to annotate instruction data. Active Instruction Tuning Kung et al. (2023) filters tasks based on prompt uncertainty, while Nuggets Li et al. (2024h) employs two-stage scoring to select diverse data. However, these data selection methods focus on filtering out low-diversity data and fall short of fundamentally enriching the instruction dataset by adding novel instructions.
@@ 2.1.2 Instruction Data Synthesis
Data synthesis methods aim to generate diverse instruction data and improve the robustness of LLMs Hu et al. (2024); Ma et al. (2022); Li et al. (2022b, 2024g, 2024e); Ye et al. (2023a); Yu et al. (2024b); Li et al. (2025b, 2024d). Some work leverages the generative capabilities of LLMs to create new instructions Taori et al. (2023); He et al. (2024); Kou et al. (2024), utilizing semantic parsing Zhao et al. (2024), transforming simple queries into complex tasks Xu et al. (2024) and blending model outputs with human-written content Chen et al. (2024b), effectively enhancing dataset diversity and quality. However, these data synthesis methods typically depend on powerful external models or extensive human annotation, leading to high computational costs and potential data leakage risks. Different from them, our MDIT is entirely labor-free, generating diverse tasks without external resources. By incorporating diversity-based clustering, we further ensure the variety of the training data.
@@ 2.2 Mixup Methods in Computer Vision
To enhance data diversity, Zhang et al. (2018) introduced Mixup for computer vision Liu et al. (2022), which creates new training samples by linearly interpolating pairs of input images and their corresponding labels. Numerous Mixup variants Yun et al. (2019); Qin et al. (2020); Kim et al. (2020); Chen et al. (2022); Wang et al. (2024); Sun et al. (2024); Islam et al. (2024) show improvements in tasks such as image classification and object detection, highlighting their effectiveness in enhancing data diversity and model robustness. However, these Mixup methods primarily focus on blending samples from similar categories within the same task. Significantly different from them, our MDIT performs interpolation across multiple tasks, generating more diverse and dynamic training data and enhancing LLM’s ability to handle complex challenges.

### v2

@@ Abstract
Yangning Li22footnotemark: 2    Zihua Lan22footnotemark: 2    Zihua Lan    Lv Qingsong     Yinghui Li    Hai-Tao Zheng33footnotemark: 3 Tsinghua University
@@ 1 Introduction
Instruction tuning has enabled large language models (LLMs) to accurately follow human instructions and significantly enhance their performance Longpre et al. (2023); Zhang et al. (2023); Yi et al. (2024). The diversity of instruction datasets plays an essential role in improving LLM’s ability to handle various scenarios Muscato et al. (2024); Fan et al. (2025). Therefore, recent research focuses on curating high-diversity and wide-ranging instruction datasets Mukherjee et al. (2023); Chung et al. (2024).
@@ 2.1.1 Instruction Data Selection
Data selection methods aim to filter out and remove low-diversity instruction data, including metric-based and model-based methods Chen et al. (2023b); Qiu et al. (2024). Metric-based selections Gonen et al. (2022); Zhou et al. (2023); Zeng et al. (2025) use quantitative metrics to identify diverse instruction data. Instruction mining Cao et al. (2023) uses a linear equation to evaluate instruction quality, while InstructionGPT-4 Wei et al. (2023) further filters multimodal instruction data Yu et al. (2024); Chen et al. (2024a) according to CLIP scores Radford et al. (2021) and instruction length. Model-based selections Wu et al. (2023); Chen et al. (2023a); Yu et al. (2023); Ge et al. (2024) leverage LLMs as data selectors to identify more diverse instructions Li et al. (2024a); Liu et al. (2024). INSTAG Lu et al. (2023) utilizes ChatGPT to annotate instruction data. Active Instruction Tuning Kung et al. (2023) filters tasks based on prompt uncertainty, while Nuggets Li et al. (2024b) employs two-stage scoring to select diverse data. However, these data selection methods focus on filtering out low-diversity data and fall short of fundamentally enriching the instruction dataset by adding novel instructions.
@@ 2.1.2 Instruction Data Synthesis
Data synthesis methods aim to generate diverse instruction data and improve the robustness of LLMs. Some work leverages the generative capabilities of LLMs to create new instructions Taori et al. (2023); He et al. (2024); Kou et al. (2024), utilizing semantic parsing Zhao et al. (2024), transforming simple queries into complex tasks Xu et al. (2024) and blending model outputs with human-written content Chen et al. (2024b), effectively enhancing dataset diversity and quality. However, these data synthesis methods typically depend on powerful external models or extensive human annotation, leading to high computational costs and potential data leakage risks. Different from them, our MDIT is entirely labor-free, generating diverse tasks without external resources. By incorporating diversity-based clustering, we further ensure the variety of the training data.
@@ 2.2 Mixup Methods in Computer Vision
To enhance data diversity, Zhang et al. (2018) introduced Mixup for computer vision, which creates new training samples by linearly interpolating pairs of input images and their corresponding labels. Numerous Mixup variants Yun et al. (2019); Qin et al. (2020); Kim et al. (2020); Chen et al. (2022); Wang et al. (2024); Sun et al. (2024); Islam et al. (2024) show improvements in tasks such as image classification and object detection, highlighting their effectiveness in enhancing data diversity and model robustness. However, these Mixup methods primarily focus on blending samples from similar categories within the same task. Significantly different from them, our MDIT performs interpolation across multiple tasks, generating more diverse and dynamic training data and enhancing LLM’s ability to handle complex challenges.